OpenStudy (anonymous):

Divide f(x) by d(x), and write a summary statement in the form indicated. f(x) = x4 + 4x3 + 6x2 + 4x + 5; d(x) = x2 + 1

hero (hero):

\(f(x) = x^4 + 4x^3 + 6x^2 + 4x + 5\); \(d(x) = x^2 + 1\) And they want the quotient in the form \(\dfrac{f(x)}{d(x)}\) right?

OpenStudy (anonymous):

Please see attached.